# NLI under the Microscope: What Atomic Hypothesis Decomposition Reveals

### v1

@@ 5.2 Critical Atoms as Questions Under Discussion (QUD)
These questions function as questions under discussion (QUD), a well-studied linguistic phenomenon Benz and Jasinskaja (2017). The update “The man smiles on his boat” has no effect on the atom about vest color and hence does not pick out the QUD associated with that atom. We call the atoms above the critical atom for each corresponding update, as they uniquely pick out particular QUDs over others. Critical atoms correspond to the particular inference or piece of knowledge a defeasible NLI example aims to test. As such, we use them as a framework to measure consistency as well as the diversity of the $\delta$-Snli dataset.
@@ 6 Measuring Inferential Consistency in $\delta$-nli
Figure 4: Grouping examples by their critical atom(s) allows us to understand under which contexts (premises and updates) a model has understood a piece of knowledge. Here, we show two $\delta$-nli examples that evaluate the same atom (top): one that strengthens it (left), and one that weakens it (right). A model that truly understands a fact and the factors that influence it (or, conversely does not) should yield consistently correct or incorrect predictions. However, some models have mixed accuracy among examples targeting the same atom, indicating that they only understand the inference under some contexts.
@@ Results.
All models exhibit room for improvement in inferential consistency (Table 4), giving us a sense of how well the LLMs we analyze have internalized the critical atomic facts evaluated in $\delta$-Snli-test. One source of inconsistency we observe arises when certain contexts (premise-update pairs) demand more implicit reasoning, multi-hop reasoning, or background knowledge. For example, consider two different contexts under which the model must evaluate the same critical atom “The people are tall”:
@@
These analyses help us understand whether models have internalized certain pieces of knowledge and the factors that influence them, raising interesting questions about how best to collect updates to increase the coverage of a fact or situation. Defeasible reasoning systems must be able to deftly modulate inferences in a manner that is sensitive to diverse contexts. Future work may leverage the identification of critical atomic sub-problems to nudge annotators toward underrepresented critical atoms or contexts.
@@ 7 Conclusion
Decomposing hypotheses of NLI and defeasible NLI problems to form atomic sub-problems allows us to measure the logical consistency of models as well as better understand the structure of non-monotonic reasoning. We find that labels of atomic sub-problems in defeasible reasoning share a more complex relation to the full problem than in traditional NLI, even within the same label class. We introduce critical atoms as the primary fact or piece of knowledge evaluated by a defeasible NLI example, enabling us to group examples by shared critical atoms and measure the inferential consistency of LLMs across different contexts.
@@ Acknowledgments
We thank the anonymous reviewers as well as the members of the University of Maryland CLIP lab for their thoughtful and thorough feedback. This work was supported by NIH Award No. R01MD016037 (Srikanth) and NSF CAREER Award No. 2339746 (Rudinger). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health or the National Science Foundation. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the paper.

### v2

@@ 5.2 Critical Atoms as Questions Under Discussion (QUD)
These questions function as questions under discussion (QUD), a well-studied linguistic phenomenon Benz and Jasinskaja (2017). The update “The man smiles on his boat” has no effect on the atom about vest color and hence does not pick out the QUD associated with that atom. We call the atoms above the critical atom for each corresponding update, as they uniquely pick out particular QUDs over others. Critical atoms correspond to the particular inference or piece of knowledge a defeasible NLI example aims to test. As such, we use them as a framework to measure consistency as well as the diversity of the $\delta$-Snli dataset. This formulation of QUDs relates to question answering-based semantics He et al. (2015); Pyatkin et al. (2021); Klein et al. (2022) in which QA pairs capture semantic information, such as semantic roles.
@@ 6 Measuring Inferential Consistency in $\delta$-nli
Figure 4: Grouping examples by their critical atom(s) allows us to understand under which contexts ($P+U$) a model has understood a piece of knowledge. Here, we show two $\delta$-nli examples that evaluate the same atom (top): one that strengthens it (left), and one that weakens it (right). A model that truly understands a fact and the factors that influence it (or, conversely does not) should yield consistently correct or incorrect predictions. However, some models have mixed accuracy among examples targeting the same atom, indicating that they only understand the inference under some contexts.
@@ Results.
All models exhibit room for improvement in inferential consistency (Table 4), giving us a sense of how well the LLMs we analyze have internalized the critical atomic facts evaluated in $\delta$-Snli-test. One source of inconsistency we observe arises when certain contexts (premise-update pairs) demand more implicit reasoning or background knowledge. For example, consider two different contexts under which the model must evaluate the same critical atom “The people are tall”:
@@
These analyses help us understand whether models understand pieces of knowledge and the factors that influence them, raising interesting questions about how best to collect updates to increase the coverage of a fact or situation. Defeasible reasoning systems must be able to deftly modulate inferences in a manner that is sensitive to diverse contexts. Future work may leverage the identification of critical atomic sub-problems to nudge annotators toward underrepresented critical atoms or contexts.
@@ 7 Conclusion
Hypothesis decomposition allows us to measure the logical consistency of models as well as better understand the structure of non-monotonic reasoning. We find that labels of atomic sub-problems in defeasible reasoning share a more complex relation to the full problem than in traditional NLI, even within the same label class. We introduce critical atoms as the primary fact evaluated by a defeasible NLI example, enabling us to group examples by shared critical atoms and measure the inferential consistency of LLMs across different contexts.
@@ Acknowledgments
We thank the anonymous reviewers as well as the members of the University of Maryland CLIP lab for their thoughtful and thorough feedback. We also thank Joe Stacey for feedback on an earlier draft of this paper. This work was supported by NIH Award No. R01MD016037 (Srikanth) and NSF CAREER Award No. 2339746 (Rudinger). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health or the National Science Foundation. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the paper.